OpenStudy (anonymous):

Which equation represents a line perpendicular to a line that passes through (-2,4) and (1,-5)

OpenStudy (anonymous):

I know how to get the slope of the line you are looking for. There are an infinite variety of equations that will be perpendicular to the line between the 2 given points. You have to provide us the answer choices. Do you know how to get the slope of the line through (-2,4) and (1,-5)?

OpenStudy (anonymous):

the answer choices are y=-3x+2 y= -1/3x +2 y= 1/3x+2 y= 3x+2

OpenStudy (anonymous):

Do you know how to get the slope of the line through (-2,4) and (1,-5)?

OpenStudy (texaschic101):

to find the slope of a line you have to use the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-2,4) x1 = -2, y1 = 4 (1,-5) x2 = 1, y2 = -5 slope(m) = (-5 - 4)/ (1 - (-2) m = -9/(1 + 2) m = -9/3 m = -3 are you understand this ?

OpenStudy (anonymous):

Is the slope -9/3? If so, then I know how to find slope lol I did m= y2-y1/x2-x1

OpenStudy (texaschic101):

the slope is -3, because -9/3 can be simplified. However, we are looking for a perpendicular line. That means the slope is going to be the negative reciprocal of -3. All that means is that you take the slope, "flip" it, and change the sign. So -3 (-3/1) will then be 1/3 for your slope. Do you understand ?

OpenStudy (anonymous):

Yes. So that means that the perpendicular slope would be -1/3x + 2?

OpenStudy (texaschic101):

lets check.....we will now use point slope formula because we have the slope and you can use either of your points. y - y1 = m(x - x1) using slope 1/3 and points (-2,4) y - 4 = 1/3(x - (-2) y - 4 = 1/3(x + 2) y - 4 = 1/3x + 2/3 y = 1/3x + 2/3 + 4 y = 1/3x + 2/3 + 12/3 y = 1/3x + 14/3 I must be doing something wrong because I am not getting any of your answer choices.

OpenStudy (texaschic101):

let me try the other points (1,-5) y - y1 = m(x - x1) y - (-5) = 1/3(x - 1) y + 5 = 1/3x - 1/3 y = 1/3x - 1/3 - 5 y = 1/3x - 5/15 - 75/15 y = 1/3x - 80/15 y = 1/3x - 16/3 I am not sure of the answer

OpenStudy (anonymous):

I have no idea, I'm doing this in a math packet

OpenStudy (texaschic101):

I am not getting any of your answer choices

OpenStudy (texaschic101):

@ivettef365 ...can you help

OpenStudy (anonymous):

The perpendicular line doesn't have to go through one of the endpoints. You're assuming it does when you construct your point slope form. It just has to have a slope of 1/3. Only answer C has a slope of 1/3. That's the answer

OpenStudy (texaschic101):

@thomaster ...please help

OpenStudy (texaschic101):

well...I suppose BangkokGarrett is right...C is the only one that has a slope of 1/3

OpenStudy (anonymous):

Thank you! :)